OpenStudy (anonymous):

angela ate at the same restraunt four times. each time she ordered a salad and left a $5 tip. she spent a total of $54. Write and solve an equation to find the cost of each salad.... please help and i appreciate all answers but only one best answer

OpenStudy (wolf1728):

She ate there 4 times so she left $20 in tips and spent a total of $54. The price of each meal is $54 = $20 + 4*salad cost $34 = 4*salad salad = 34/4 salad = $8.50

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."